OpenStudy (anonymous):

Help with Logarithms that have uneven bases? 6^x = 24 8^(x - 2) = 3 5^(x + 5) = 9^x 3^(x - 4) = 7^(x + 9) I have a test I need to take on these tomorrow and I'm REALLY freaking over because I just don't understand them ); If you could take me through the steps of each ones (or just go through one and provide an answer key so I can check my answers against it) that would be awesome! Thanks!

OpenStudy (anonymous):

\[ 6^x=24\\ x=\log_624\\ x=\frac{\ln24}{\ln6} \] \[ 8^{x-2}=3\\ x-2=\log_83 \]Do you follow?

OpenStudy (anonymous):

I think so ); But ... Gah, probably not. I'm so incredibly bad at these things. Sorry for taking up your time like this! So, with 6^x = 24, you just switch it to a log with a base of 6 and keep 24 at the end like that? So, if it were 3^x = 5, you'd just put log3 5?

OpenStudy (anonymous):

Depends on what your teacher wants, you can simplify it a little bit more if you really need to but the answer is irrational so no matter what it'll have some \(\ln\) terms. You just remember what logarithm means, \(\log_ab=x\) is the exact same thing as saying \(a^x=b\).